how many electrons passes through a 20 ohm resistor through 10 mins if there is a drop of 30 volts a cross it


1
Expert's answer
2021-09-07T14:07:17-0400

From Ohm’s Law

"V=IR \\

I=\frac{V}{R}= \frac{30 \;V}{20} = 1.5 \;A \\

I= \frac{Q}{t} \\

Q=It = (1.5 \;A) \times (10 \times 60 \;s) = 900 \;C \\

Q=ne \\

n= \frac{Q}{e} = \frac{900 \;C}{1.6 \times 10^{-19} \;C}=562.5 \times 10^{19}"